# Unified quantification of entanglement and magic in information scrambling and their trade-off relationship

### v1

@@ Abstract
Information scrambling is a phenomenon observed in a wide range of quantum systems, in which initially localized information becomes distributed across the entire system. Clarifying the essence of this phenomenon is made possible by quantum resource theory, which provides an axiomatic framework for quantifying various forms of quantumness. Within this framework, information scrambling can be classified into two distinct types: entanglement scrambling and magic scrambling. Although each has been analyzed individually through their respective resource theories, such separate treatments fail to capture the inherent trade-off between them. In this study, we introduce a unified measure of the two types of scrambling, grounded in resource-theoretic principles. This unified approach reveals a rigorous trade-off relationship between entanglement and magic scrambling, as the exact maximum value of the proposed measure can be derived analytically. Furthermore, we quantify the scrambling capability of unitary transformations in terms of their ability to amplify this measure.
I Introduction
@@
Quantum systems exhibit a range of non-classical and counterintuitive behaviors. Understanding the origins of such quantumness remains one of the most fundamental challenges in quantum mechanics. Quantum resource theories provide a powerful framework to capture and formalize the essential features of quantum systems (see Ref. [1] for a review). In a typical resource-theoretic framework, a fixed set of operations—known as free operations—is specified in advance, and any property that cannot be generated using only these operations is regarded as a resource. To quantify the resource, a resource measure (a function that assigns a numerical value to the resource) is introduced, typically satisfying several core properties such as faithfulness and monotonicity. This formalism was originally developed to quantify entanglement. Since then, a variety of quantum features, including coherence [2, 3, 4, 5], superposition [6], magic [7], uncomplexity [8], and thermodynamic nonequilibrium [9, 10], have been studied within this framework. Constructing a resource theory for a given quantum property enables both its characterization and quantitative analysis.
Chaos has been extensively studied in mathematics and physics. One of the defining features of chaos is the butterfly effect, in which a slight change in initial conditions can lead to drastically different outcomes. In physical systems, chaotic behavior is observed in many-body systems with complex interactions, as well as in black holes. In quantum systems, chaos gives rise to quantum information scrambling, a process in which information initially localized in a small subsystem becomes distributed throughout the entire system. This scrambling has recently been linked to a wide range of fields, including quantum error correction [11], quantum machine learning [12, 13, 14, 15], or Clifford decoder [16]. Viewing information scrambling as a resource provides a powerful framework for understanding and quantifying this phenomenon.
Information scrambling has been quantified using various measures, such as the out-of-time-ordered correlator (OTOC) [17, 18, 19], the entanglement entropy [20, 21], and the stabilizer Rényi entropy [22, 23, 24]. Since these measures capture different aspects of scrambling, ambiguity remains as to what constitutes a highly scrambling transformation. Recently, a rigorous framework for the resource theory of information scrambling has been proposed [25, 26]. In the resource-theoretic context, scrambling is categorized into two distinct types based on underlying mechanisms: entanglement scrambling and magic scrambling (see FIG. 1). While these names are not yet standardized, they suggest that each type of scrambling can be understood as a quantum resource. Entanglement scrambling refers to transformations that increase the number of non-identity components in an operator’s Pauli basis expansion. It can be quantified by measures such as Pauli growth [25] and circuit sensitivity [26], which behave as valid resource monotones when non-entangling unitaries (i.e., single-qubit unitaries and swap operations) are taken as free operations. In contrast, magic scrambling corresponds to transformations that increase the number of terms required in the Pauli basis expansion. It is quantified by measures such as OTOC magic [25] and magic power [26], which are valid resource monotones under Clifford unitaries as the set of free operation. These resource measures satisfy several fundamental properties that justify their interpretation as resource monotones. Grounded in this rigorous framework, they characterize each type of scrambling individually.
@@
In this study, we reconstruct the resource-theoretic framework for information scrambling. We define measures that treat both entanglement scrambling and magic scrambling in parallel. Furthermore, we focus on the fact that entanglement scrambling and magic scrambling cannot be simultaneously maximized. To clarify the trade-off relationship between them, we introduce a new quantity called operator complexity, which captures both types of scrambling in a unified way. As our most significant result, we analytically derive the exact maximum value of the operator complexity—revealing a fundamental limitation inherent to information scrambling. In addition, as a measure of the ability to amplify the operator complexity, we introduce a unified scrambling measure. These quantities satisfy several essential properties as resource monotones, thereby ensuring the soundness of our framework. The proofs for most propositions are provided in the appendix.
@@
II Scrambling of operators
Information scrambling can be defined as the ability of unitary transformations to amplify the complexity of operators. Entanglement scrambling and magic scrambling, which are different types of scrambling arising from distinct mechanisms, can be quantified using different measures of the complexity of operators. In this section, we introduce two measures of operators that correspond to each scrambling in the context of resource theory. Furthermore, we define a unified measure that takes into account both scrambling and derive its rigorous maximum value. This clarifies a trade-off between the two types of scrambling.
@@ II Scrambling of operators
Here, we take the base of the logarithm as a parameter $a$. If $a$ is omitted, we set $a=4$ so that the maximum value of the quantum Fourier entropy becomes $n$. This choice aligns the ranges of $W(O)$ and $S(O)$, enabling us to treat the two types of scrambling on equal footing. For any normalized operator $O$, the quantum Fourier entropy satisfies the following properties:
@@
We provide the solution to this problem in the appendix. As a result, the closed-form expression of the operator that maximizes $R_{a}(O)$ is derived as
@@
See the appendix for the details of the calculation. In particular, when the parameter $a$ is set to 4, we obtain
@@
where $\lambda=\log_{a}4$. The quantity $(1+\lambda)n$ is a trivial upper bound, since $R_{a}(O)=W(O)+\lambda S(O)$. Eq. (10) also captures the fundamental limitation inherent in the two types of scrambling.
In FIG. 2, we map the several representative operators on the $W$-$S$ plane. First, any weight-$n$ Pauli operator—for example, $O_{1}=Z_{1}\otimes\cdots\otimes Z_{n}$—gives the maximum value of $W(O)$, while $S_{a}(O)$ remains zero. Linear combinations of weight-$n$ Pauli operators also attain the maximum value of $W(O)$, and although they yield nonzero values of $S_{a}(O)$, they do not maximize it. For instance, the operator $O_{2}$ shown in FIG. 2 maximizes $W(O)$ and achieves $S_{a}(O_{2})=n\log_{a}3$, which is smaller than the maximum value of $S_{a}(O)$. In contrast, the uniform linear combination of all Pauli basis elements with equal amplitude, $O_{3}=1/2^{n}\sum_{i}P_{i}$, maximizes $S_{a}(O)$, while its average Pauli weight is not maximized: $W(O_{3})=3n/4$. Most importantly, $O_{\mathrm{Max}}$ is optimal with respect to the operator complexity $R_{a}(O)$, even though it does not individually maximize $W(O)$ or $S_{a}(O)$. The exact location of $O_{\mathrm{Max}}$ as a function of the parameter $a$ is also illustrated in FIG 2. In the limit $a\rightarrow 1$, $O_{\mathrm{Max}}$ converges to $O_{3}$, whereas in the limit $a\rightarrow\infty$, it converges to $O_{2}$.
Figure 2: Overview of the $W$-$S$ plane for arbitrary linear operators. A key observation is the trade-off relation between the two types of scrambling, illustrated by the solid red curve. When $W(O)$ reaches its maximum, $S_{a}(O)$ cannot, and vice versa—when $S_{a}(O)$ is maximized, $W(O)$ takes a specific, non-maximal value. The dashed curve provides a rough sketch of the general trend (possibly imprecise).
III Scrambling of unitaries
The measures introduced in the previous section allow us to quantify information scrambling exhibited by unitary transformations. The average Pauli weight and the quantum Fourier entropy are associated with entanglement scrambling and magic scrambling, respectively. Furthermore, the operator complexity provides a unified framework for quantifying information scrambling.
@@ III Scrambling of unitaries
These support the validity of $\tilde{G}(U)$ as a resource monotone. In particular, maxitivity—a stronger property than subadditivity—is a distinctive feature of $\tilde{G}(U)$. $\tilde{G}(U)$ can be regarded as a more suitable measure for quantifying the information scrambling of a localized unitary. For instance, consider a localized unitary of the form $U=U_{1}\otimes I_{2}$. In this case, $\tilde{G}(U)=\tilde{G}(U_{1})$ can accurately evaluate the scrambling capability, whereas $G(U)\leq G(U_{1})$ may dilute it. In addition, since the set of operators satisfying $W(O)=1$ and $S(O)=0$ coincides with the set of weight-1 Pauli operators, $\tilde{G}(U)$ has the advantage of being numerically computable. Taken together, these observations suggest that $\tilde{G}(U)$ is a promising candidate for a unified measure of information scrambling.
  Operator measure   Unitary measure   Free unitary
IV Conclusion
In this study, we constructed a unified framework of resource theory for information scrambling. In particular, we introduced the concept of operator complexity, which accounts both entanglement scrambling and magic scrambling. As a main result, we derived the exact maximum value of this measure, revealing a trade-off relationship inherent in information scrambling. In addition, we quantified the scrambling capability intrinsic to unitary transformations as the amplification of such measures. For the complexity growth, we also introduced an alternative definition, which serves as another plausible resource measure. These measures exhibit a parallel structure and satisfy the properties required of resource monotones.
We anticipate that this work will contribute to a deeper understanding of the essence of information scrambling. Moreover, since our framework is built on measures for general linear operators, it is expected to extend naturally to more general settings, such as non-unitary channels. We hope that this study will inspire further developments across a broad range topics in quantum physics.
@@ Acknowledgements.
This appendix provides detailed proofs of the propositions presented in the main text.
Appendix A Average Pauli weight
We present the basic properties of the average Pauli weight $W(O)$ and provide reconstructed proofs based on Ref. [25].
Proof: Faithfulness.
By definition of the average Pauli weight, $W(O)\geq 0$ holds trivially. Equality holds if and only if $O$ is the identity operator. If $O$ is traceless, it contains no identity term in its Pauli expansion. Thus, for any traceless operator, $W(O)\geq 1$ holds, and $W(O)=1$ if and only if the Pauli expansion of $O$ consists solely of weight-1 Pauli operators. ∎
Proof: Invariance..
We prove the invariance of the average Pauli weight: for nay operator $O$, if $U$ is a non-entangling unitary, then $W(U^{\dagger}OU)=W(O)$. Here, a non-entangling unitary refers to any unitary that can be generated by single-qubit unitaries and swap operations.
In this proof, we expand the operator as $O=\sum_{\vec{a}}c_{\vec{a}}P_{\vec{a}}$, where the index $j$ denotes the position of the local Pauli operator. The average Pauli weight can then be computed as follows:
$W(O)$ $=\sum_{j}\sum_{\vec{a}(P_{j}\neq I)}|c_{\vec{a}}|^{2}$
$=\sum_{j}\frac{1}{d^{2n}}\sum_{\vec{a}(P_{j}\neq I)}\left|\mathrm{Tr}[OP_{\vec{a}}]\right|^{2}$
$=\sum_{j}\left[\frac{1}{d^{2n}}\sum_{\vec{a}}\left|\mathrm{Tr}[OP_{\vec{a}}]\right|^{2}-\frac{1}{d^{2n}}\sum_{\vec{a}(P_{j}=I)}\left|\mathrm{Tr}[OP_{\vec{a}}]\right|^{2}\right]$
$=\sum_{j}\left[\frac{1}{d^{n}}\mathrm{Tr}[O^{\dagger}O]-\frac{1}{d^{n+1}}\mathrm{Tr}\left[(\mathrm{Tr}_{j}[O])^{\dagger}(\mathrm{Tr}_{j}[O])\right]\right].$ (15)
Here, $d=2$ and $\mathrm{Tr}_{j}$ denotes the partial trace over the $j$th qubit. We note that the partial trace operation removes all terms except those corresponding to Pauli strings with $P_{j}=I$. When $U$ is generated as a product of single-qubit unitaries, the same calculation yields the following:
$W(U^{\dagger}OU)$ $=\sum_{j}\left[\frac{1}{d^{n}}\mathrm{Tr}[(U^{\dagger}OU)^{\dagger}(U^{\dagger}OU)]-\frac{1}{d^{n+1}}\mathrm{Tr}\left[(\mathrm{Tr}_{j}[U^{\dagger}OU])^{\dagger}(\mathrm{Tr}_{j}[U^{\dagger}OU])\right]\right]$
$=\sum_{j}\left[\frac{1}{d^{n}}\mathrm{Tr}[O^{\dagger}O]-\frac{1}{d^{n+1}}\mathrm{Tr}\left[(\mathrm{Tr}_{j}[O])^{\dagger}(\mathrm{Tr}_{j}[O])\right]\right].$ (16)
The second equality follows from the fact that any local unitary operation on the $j$th qubit preserves terms with $P_{j}=I$ and cannot generate such terms from those with $P_{j}\neq I$. Thus, if $U$ can be written as a product of local unitary operations as $U=\otimes_{j}U_{j}$, then $W(U^{\dagger}OU)=W(O)$ holds. In addition, since the Pauli weight is clearly preserved under swap operations, $W(O)$ is invariant under swaps as well. Therefore, the average Pauli weight is invariant under any non-entangling unitary, which can be generated as a product of single-qubit unitaries and swap operations. ∎
Proof: Additivity..
Additivity can be proved directly. By definition, the weight of Pauli strings is additive: $W(P_{1}\otimes P_{2})=W(P_{1})+W(P_{2})$. Therefore, for arbitrary normalized operators $O_{1}=\sum_{i}c_{i}P_{i}$ and $O_{2}=\sum_{j}c_{j}^{\prime}P_{j}^{\prime}$, additivity follows from the following calculation:
$W(O_{1}\otimes O_{2})$ $=\sum_{i,j}|c_{i}c_{j}^{\prime}|^{2}W(P_{i}\otimes P_{j}^{\prime})$
$=\sum_{i}|c_{i}|^{2}W(P_{i})+\sum_{j}|c_{j}^{\prime}|^{2}W(P_{j}^{\prime})$
$=W(O_{1})+W(O_{2}).$ (17)
Appendix B Quantum Fourier entropy
We provide proofs of the basic properties of the quantum Fourier entropy $S_{a}(O)$.
Proof: Faithfulness..
Faithfulness follows immediately from the definition of $S_{a}(O)$, namely, its range is from 0 to $\log_{a}(4^{n})$. In particular, when the parameter $a$ is set to 4, the maximum value of $S(O)$ is $n$. The equality $S_{a}(O)=0$ holds if and only if $O$ is a Pauli operator, since this corresponds to the case where the probability distribution is concentrated on a single element. ∎
Proof: Invariance..
Invariance follows directly from the definition of Clifford unitaries. If $U$ is a Clifford unitary and $P$ is a Pauli operator, then $U^{\dagger}PU$ is also a Pauli operator, and the mapping is injective. Therefore, when $O=\sum_{i}c_{i}P_{i}$ is conjugated by a Clifford unitary, it transforms into $U^{\dagger}OU=\sum_{i}c_{i}^{\prime}P_{i}$, where $\{c_{i}^{\prime}\}_{i}$ is a permutation of the original coefficients $\{c_{i}\}_{i}$. Since the value of the quantum Fourier entropy is invariant under permutations of the coefficients, it remains unchanged under any Clifford unitary transformation. ∎
Proof: Additivity..
Additivity can be derived through a straightforward argument. For normalized operators $O_{1}=\sum_{i}c_{i}P_{i}$ and $O_{2}=\sum_{j}c^{\prime}_{j}P^{\prime}_{j}$, their tensor product is expressed as $O_{1}\otimes O_{2}=\sum_{i,j}c_{i}c^{\prime}_{j}P_{i}\otimes P^{\prime}_{j}$. Here, $\{P_{i}\otimes P^{\prime}_{j}\}_{ij}$ forms the Pauli group on the total Hilbert space, and the coefficients $\{c_{i}c^{\prime}_{j}\}_{ij}$ satisfy the normalization condition $\sum_{ij}|c_{i}c^{\prime}_{j}|^{2}=1$. Therefore, we can obtain the following equality:
$S_{a}(O_{1}\otimes O_{2})$ $=-\sum_{i,j}|c_{i}c^{\prime}_{j}|^{2}\log|c_{i}c^{\prime}_{j}|^{2}$
$=-\sum_{i}|c_{i}|^{2}\log|c_{i}|^{2}-\sum_{j}|c^{\prime}_{j}|^{2}\log|c^{\prime}_{j}|^{2}$
$=S_{a}(O_{1})+S_{a}(O_{2}).$ (18)
Appendix C Operator complexity

### v2

@@ Abstract
Entanglement and magic are among the most fundamental properties unique to quantum systems. Each quantity captures a different aspect of non-classical behavior, and each can be regarded as a resource within its own operational setting. However, the interrelation between them has not yet been fully clarified, and whether a more fundamental measure exists remains an open question. Addressing these issues is essential for deepening our understanding of quantumness. In this study, we establish a unified resource theory of information scrambling, consisting of two types: entanglement scrambling and magic scrambling. We introduce a measure that jointly characterizes both types of scrambling. This unified approach reveals a rigorous trade-off relation between entanglement and magic scrambling, as the exact maximum value of the proposed measure can be derived analytically. Furthermore, we quantify the scrambling capability of unitary transformations in terms of their ability to amplify this measure. Our work provides insights into the connection between entanglement and magic that extend beyond the context of information scrambling.
@@
Introduction.—Quantum systems exhibit a variety of non-classical and counterintuitive behaviors. Understanding the origins of such quantumness remains one of the most fundamental challenges in quantum mechanics. Quantum resource theories [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] provide a powerful framework to capture and formalize the essential features of quantum systems (see Ref. [26] for a review). In a typical framework, a fixed set of operations, known as free operations, is specified in advance, and any property that cannot be generated using only these operations is regarded as a resource. Constructing a resource theory for a given quantum property enables both its characterization and quantitative analysis.
Entanglement and magic are key characteristics unique to quantum systems. Each of these properties embodies quantumness, but they characterize different aspects, and their relationship remains unclear in several respects. For instance, the GHZ state is highly entangled and exhibits strong quantum behavior. However, it provides no quantum advantage from the perspective of stabilizer computation, as it can be generated solely by Clifford unitaries. Conversely, the concatenation of T states exhibits large magic and enables quantum computational supremacy, even though it contains no entanglement. From these observations, one natural question arises: is there any other crucial quantity that represents quantumness? Recent studies have explored the overlap between entanglement and magic [27, 28, 29, 30, 31]. Importantly, neither resource on its own can fully account for the source of quantum advantage in certain contexts, such as stabilizer tensor networks [32]. The existence of a unified resource measure that incorporates both resources has been suggested, although its explicit form has not yet been specified [32, 33].
@@
In this study, we establish a unified resource-theoretic framework of information scrambling and derive the rigorous trade-off relation between entanglement and magic within this framework. Information scrambling refers to the phenomenon in which information initially localized in a small subsystem becomes distributed throughout the entire system [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. In the resource-theoretic context, scrambling is categorized into two distinct types based on their underlying mechanisms [47, 48]: entanglement scrambling and magic scrambling (see FIG. 1). Each type can be regarded as a resource under its own set of free operations. However, treating them separately fails to reveal how they may influence or complement one another. We define measures that treat both entanglement scrambling and magic scrambling in parallel. Furthermore, we emphasize that entanglement scrambling and magic scrambling cannot be simultaneously maximized. To clarify their interrelationship, we introduce a measure, termed operator complexity, which captures both types of scrambling in a unified manner. As our most important result, we analytically derive the exact maximum value of the operator complexity, thereby revealing a fundamental trade-off inherent to information scrambling. This discovery is expected to advance our understanding of how entanglement and magic are connected beyond the scope of information scrambling. In addition, we propose a unified scrambling measure as an indicator of the ability to amplify the operator complexity. These quantities satisfy several essential properties of resource monotones, ensuring the consistency and soundness of our framework. Proofs for most propositions are provided in [49] and the End Matter.
@@
Scrambling of operators.—Information scrambling can be defined as the ability of unitary transformations to amplify the complexity of operators. Entanglement scrambling and magic scrambling, which are different types of scrambling arising from distinct mechanisms, can be quantified using different measures of the complexity of operators. In this section, we introduce two measures of operators that correspond to each scrambling in the context of resource theory. Furthermore, we define a unified measure that takes into account both scrambling and derive its rigorous maximum value. This clarifies the fundamental trade-off between the two types of scrambling.
@@
The advantages of this operator-based measure are discussed in Ref. [50]. Here, we take the base of the logarithm as a parameter $a$. If $a$ is omitted, we set $a=4$ so that the maximum value of the quantum Fourier entropy becomes $n$. This choice aligns the ranges of $W(O)$ and $S(O)$, enabling us to treat the two types of scrambling on equal footing. For any normalized operator $O$, the quantum Fourier entropy satisfies the following properties:
@@
We provide the solution to this problem in the End Matter. As a result, the closed-form expression of the operator that maximizes $R_{a}(O)$ is derived as
@@
See the End Matter for the details of the calculation. In particular, when the parameter $a$ is set to 4, we obtain
@@
where $\lambda=\log_{a}4$. The quantity $(1+\lambda)n$ is a trivial upper bound, since $R_{a}(O)=W(O)+\lambda S(O)$. Eq. (10) also captures the intrinsic limitation between the two types of scrambling.
Figure 2: Overview of the $W$-$S$ plane for arbitrary linear operators. A key observation is the trade-off relation between the two types of scrambling, illustrated by the solid red curve. When $W(O)$ reaches its maximum, $S_{a}(O)$ cannot, and vice versa: when $S_{a}(O)$ is maximized, $W(O)$ takes a specific, non-maximal value. The dashed curve provides a rough sketch of the general trend (possibly imprecise).
In FIG. 2, we map the several representative operators on the $W$-$S$ plane. First, any weight-$n$ Pauli operator, such as $O_{1}=Z_{1}\otimes\cdots\otimes Z_{n}$, gives the maximum value of $W(O)$, while $S_{a}(O)$ remains zero. Linear combinations of weight-$n$ Pauli operators also attain the maximum value of $W(O)$, and although they yield nonzero values of $S_{a}(O)$, they do not maximize it. For instance, the operator $O_{2}$ shown in FIG. 2 maximizes $W(O)$ and achieves $S_{a}(O_{2})=n\log_{a}3$, which is smaller than the maximum value of $S_{a}(O)$. In contrast, the uniform linear combination of all Pauli basis elements with equal amplitude, $O_{3}=1/2^{n}\sum_{i}P_{i}$, maximizes $S_{a}(O)$, while its average Pauli weight is not maximized: $W(O_{3})=3n/4$. Most importantly, $O_{\mathrm{Max}}$ is optimal with respect to the operator complexity $R_{a}(O)$, even though it does not individually maximize $W(O)$ or $S_{a}(O)$. The exact location of $O_{\mathrm{Max}}$ as a function of the parameter $a$ is also illustrated in FIG 2. In the limit $a\rightarrow 1$, $O_{\mathrm{Max}}$ converges to $O_{3}$, whereas in the limit $a\rightarrow\infty$, it converges to $O_{2}$.
  Operator   Unitary   Free unitary
Scrambling of unitaries.—The measures introduced in the previous section allow us to quantify information scrambling exhibited by unitary transformations. The average Pauli weight and the quantum Fourier entropy are associated with entanglement scrambling and magic scrambling, respectively. Furthermore, the operator complexity provides a unified framework for quantifying information scrambling.
@@
These support the validity of $\tilde{G}(U)$ as a resource monotone. In particular, maxitivity, which is a stronger property than subadditivity, is a distinctive feature of $\tilde{G}(U)$. $\tilde{G}(U)$ can be regarded as a more suitable measure for quantifying the information scrambling of a localized unitary. For instance, consider a localized unitary of the form $U=U_{1}\otimes I_{2}$. In this case, $\tilde{G}(U)=\tilde{G}(U_{1})$ can accurately evaluate the scrambling capability, whereas $G(U)\leq G(U_{1})$ may dilute it. In addition, since the set of operators satisfying $W(O)=1$ and $S(O)=0$ coincides with the set of weight-1 Pauli operators, $\tilde{G}(U)$ has the advantage of being numerically computable. Taken together, these observations suggest that $\tilde{G}(U)$ is a promising candidate for a unified measure of information scrambling.
Conclusion.—In this study, we constructed a unified framework of resource theory for information scrambling. In particular, we introduced the concept of operator complexity, which accounts both entanglement scrambling and magic scrambling. As a main result, we derived the exact maximum value of this measure, revealing a trade-off relation inherent in information scrambling. In addition, we quantified the scrambling capability intrinsic to unitary transformations as the amplification of such measures. For the complexity growth, we also introduced an alternative definition, which serves as another plausible resource measure. These measures exhibit a parallel structure and satisfy the properties required of resource monotones.
As for future directions, developing efficient methods to compute concrete values of scrambling measures will be an important task. Furthermore, our framework can likely be extended to non-unitary channels, allowing the quantification of their scrambling capability. We expect that this work will deepen the understanding of the fundamental nature of information scrambling. Moreover, our analysis offers perspectives on the connection between entanglement and magic that go beyond the scope of scrambling itself. We hope that this study will inspire further developments across a broad range of topics in quantum physics.
@@ Acknowledgements.
Appendix A Operator complexity